# ESG-Net: Event-Aware Semantic Guided Network for Dense Audio-Visual Event Localization

### v1

@@ Abstract
Dense audio-visual event localization (DAVE) aims to identify event categories and locate the temporal boundaries in untrimmed videos. Most studies only employ event-related semantic constraints on the final outputs, lacking cross-modal semantic bridging in intermediate layers. This causes modality semantic gap for further fusion, making it difficult to distinguish between event-related content and irrelevant background content. Moreover, they rarely consider the correlations between events, which limits the model to infer concurrent events among complex scenarios. In this paper, we incorporate multi-stage semantic guidance and multi-event relationship modeling, which respectively enable hierarchical semantic understanding of audio-visual events and adaptive extraction of event dependencies, thereby better focusing on event-related information. Specifically, our event-aware semantic guided network (ESG-Net) includes a early semantics interaction (ESI) module and a mixture of dependency experts (MoDE) module. ESI applys multi-stage semantic guidance to explicitly constrain the model in learning semantic information through multi-modal early fusion and several classification loss functions, ensuring hierarchical understanding of event-related content. MoDE promotes the extraction of multi-event dependencies through multiple serial mixture of experts with adaptive weight allocation. Extensive experiments demonstrate that our method significantly surpasses the state-of-the-art methods, while greatly reducing parameters and computational load. Our code will be released on https://github.com/uchiha99999/ESG-Net.
Index Terms:
Audio-visual event localization, cross-modal semantic bridging, multi-modal early fusion, mixture of experts.
@@ I Introduction
In recent years, most existing cross-modal frameworks [1, 5] for detecting dense events only perform event-related constraints on the final outputs. However, there exists a deficiency in cross-modal semantic bridging in intermediate layers, leading to modality semantic gap for further multi-modal fusion. Recent methods [6] attempt to use contrastive learning [7] at the early stages of the model for cross-modal alignment, thereby enhancing the correlations of different modalities. But forcibly aligning multi-modal features without semantic constraints may introduce additional noise due to mismatched segments within untrimmed videos (e.g., a scene of playing tennis accompanied by the sound of cheers at the last frame in Figure 1). In summary, most existing methods fail to impose sufficient cross-modal semantic bridging in intermediate layers, making it difficult for the model to distinguish between event-related content and irrelevant background content.
@@
Figure 1: (a) Most previous methods only performe semantic constraints on the final outputs. (b) We consider event-related semantic bridging between different modalities in intermediate layers through cross-modal early fusion and semantic guidance.
@@
Additionally, few studies take the correlations between audio-visual events into consideration, which limits the model’s ability to infer concurrent events in complex scenarios. Specifically, there are complex dependencies between different audio-visual events [1] (e.g., thunder is often accompanied by rain). These dependencies are variable (e.g., piano solo or playing together with guitar) and may differ in order and overlap duration. How to adaptively capture those diverse dependencies in complex scenarios to improve the reasoning capability of model for concurrent events is also a crucial issue for DAVE.
@@
In this work, we propose an event-aware semantic guided network (ESG-Net), which leverages multi-stage semantic guidance for hierarchical audio-visual events understanding and multi-event relationship modeling for adaptive dependencies extraction, respectively. (1) As shown in Figure 1, compared with previous methods, our model add multi-modal early fusion to prepare for cross-modal semantic bridging in intermediate layers (audio-visual events must be visible and audible simultaneously). Then, we perform event-related semantic alignment in intermediate layers to reduce the modality semantic gap and guide the model to focus on foreground information (e.g., audio-visual events like "playing tennis" tagged in Figure 1). (2) Additionally, we also consider the dependencies among multiple events and capture these complex multi-event relationships by mixture of experts (MoE) [8, 9, 10] modules. Overall, our method hierarchically emphasizes event-related content through multi-stage processing, where the semantic guidance and dependency modeling jointly enable accurate audio-visual event detection.
@@
Specifically, ESG-Net consists of an early semantics interaction (ESI) module and a mixture of dependency experts (MoDE) module. (1) The ESI module includes multi-modal early fusion and multi-stage semantic guidance. Firstly, we fuse cross-modal information gradually through three stages: single-modal attention, audio(visual)-driven mixture, cross-modal pyramid. Then, we perform cross-modal semantic bridging in each stage through weighted classification loss functions, which enable event-related sementic guidance for our model. (2) MoDE is proposed to address the the modeling of correlations between audio-visual events with efficient computation. MoDE consists of multiple sequential mixture of experts [8, 9, 10] layers, with each layer selecting only one expert at a time. Through adaptive selection of experts in different layers, diverse event dependencies can be captured. In summary, our ESI and MoDE sequentially achieve multi-stage semantic bridging and multi-event relationships extraction, ensuring the attention to event-related content for untrimmed videos.
@@
Our event-aware semantic guided network can gradually focus on audio-visual content across multiple intermediate layers and capture multi-event relationships for accurate dense events localization. The results show that ESG-Net surpasses the baseline model by 2.1% on I3D+VGGish [11, 12] backbone and 3.7% on ONE-PEACE [13] backbone. Our contributions are summarized as:
@@
  • We introduce an early semantics interaction module and a mixture of dependency experts module, designed for cross-modal bridging by explicit semantic guidance in intermediate layers and adaptive multi-event dependencies extraction, respectively.
  • Abundant experiments demonstrate that our event-aware semantic guided network outperforms existing state-of-the-art methods on untrimmed videos across various intersection-over-union thresholds.
@@ II-A Audio-Visual Event Localization
DAVE aims to detect events that can be seen and heard simultaneously in video clips. Tian et al. [14] first introduced the audio visual event localization (AVEL) task and AVE dataset, and proposed a dual multi-modal residual network for this task. Then some studies [15, 16] provided supervised cross-modal fusion strategies and event detection methods, while others [17, 18, 19, 20] are compatible with both weakly-supervised and supervised methods. However, these studies are based on trimmed video clips, with a single event in each clip, which are not fit with the real-world scenarios. To address this issues, Geng et al. [1] released UnAV-100, a real-world dataset containing a large number of untrimmed videos with dense events. Subsequently, they proposed a unified framework to integrate three tasks [5], including temporal action localization (TAL) [21, 22, 23], sound event detection (SED) [24, 25], and AVEL [14, 15, 26]. Afterwards, Faegheh et al. [27] improved the weakly-supervised audio-visual video parsing [28, 29, 30, 31] by employing a reference module and an anchor module to extract single-modal and multi-modal information, which resulted in better performance on untrimmed videos. Recently, Tang et al. [32] upgraded Vicuna to AVicuna, enhancing the multi-modal understanding capabilities of large language models by aligning audio and video features with text. Xing et al. [6] proposed a locality aware cross-modal correspondence learning framework and improved the inaccurate recognition of event boundaries [1] by using contrastive learning and dynamic local temporal attention. While Zhou et al. [33] improve the performance through adaptive cross-modal attention and temporal modeling at different granularities. However, these methods only employ event-related constraints on the final outputs, lacking cross-modal semantic bridging in intermediate layers, which remain modality semantic gap for insufficient multi-modal fusion. In this work, we perform multi-stage semantic guidance for multi-modal alignment, which allows the model to focus on event-related content in multiple intermediate layers, thereby improving the accuracy of event detection.
@@ II-B Audio-Visual Bridging
The joint analysis of audio and video is a powerful tool that can be applied to various tasks [34]. How to reasonably bridge different modalities is crucial for models to understand task-related content. Early researches [35, 36, 37] simply concatenated and projected the audio and visual features extracted from the sub-networks. This may result in insufficient features fusion and suboptimal extraction of task-specific information. Subsequently, researchers used the audio-visual similarity to guide attention toward relevant regions for particular tasks [14, 38, 39, 15, 40, 41]. Some methods also applied contrastive learning for self-supervised synchronous tasks [42, 43, 44]. Currently, there are typical frameworks mainly include single-modal modeling and multi-modal cross-attention, which are used for audio-visual correspondence (AVC) [45], audio-visual parsing (AVP) [46], and AVEL [18, 17, 30, 1, 47]. Most of these methods only empoly late multi-modal fusion and perform semantic constraints on the final outputs, ignoring semantic cross-modal bridging in intermediate layers, which may make the model difficult to focus on task-related content as early as possible.
Other studies attempt to use early cross-modal bridging to reduce modality gap. Introducing contrastive learning [7] is a reasonable method to strengthen the connections between audio and video. For example, Hong et al. [48] proposed hyperbolic alignment to implement zero-shot learning. Trevine et al. [49] achieved video deepfake detection by aligning the multi-modal features of the same frame. Lee et al. [50] conducted pairwise contrastive learning to map audio, images, and text into a multi-modal shared latent space, while Mo et al. [51] built a framework for localization, separation, and recognition by matching positive samples. Although these methods can effectively optimize multi-modal features, high correlation between corresponding frames is required by the original audio and video. In most untrimmed videos, there are many mismatched segments. Forcibly performing frame-level alignment may introduce significant noise. Some studies used co-attention to capture the relationships between multi-modal inputs [52, 16, 19, 20, 1]. Other methods also added a parallel early fusion branch in addition to two single-modal branches [53, 54, 55] for temporal alignment. However, these audio-visual bridging methods ignore the semantic alignment between different modalities in intermediate processes. Our method employs a multi-stage guided classification to enforce explicit cross-modal semantic bridging gradually, thereby reducing modality semantic gap and promoting the attention of our model to event-related content.
@@ III Method
To overcome the multi-modal semantic gap in intermediate layers and the modeling of event correlations for DAVE, we design an early semantics interaction module to achieve multi-stage cross-modal semantic bridging, and a mixture of dependency experts module to capture complex relationships between events. An overview of the event-aware semantic guided network we propose is shown in Figure 2.
@@ III-B Overall Architecture
ESG-Net comprises four key components: 1) pre-trained audio encoder and visual encoder to extract features from original videos; 2) an early semantics interaction module for cross-modal semantic bridging; 3) a mixture of dependency experts module for capturing multi-event dependencies; 4) a decoder for localizing audio-visual events. The five key components are executed in sequence as illustrated in Figure 2.
@@
Early Semantics Interaction Module. To achieve cross-modal semantic bridging in intermediate layers, ESI consists of Multi-Modal Early Fusion and Multi-Stage Semantic Guidance. The Multi-Modal Early Fusion includes Single-Modal Attention, Audio(Visual)-Driven Mixture and Cross-Modal Pyramid Module[1]. The outputs of these three parts are $\{\hat{F}_{A},\hat{F}_{V}\}\in\mathbb{R}^{T_{m}\times D}$, $\{F_{AV},F_{VA}\}\in\mathbb{R}^{T_{m}\times D}$, $\{\hat{F}_{AV},\hat{F}_{VA}\}\in\mathbb{R}^{T_{l}\times D}$, respectively, where $T_{m}$ is the maximum length among all videos, $T_{l}$ is the total length of the multi-level pyramid features. Then, they are used for Multi-Stage Semantic Guidance through several classfication loss functions. The details are as shown in Section III-C.
@@
Mixture of Dependency Experts Module. After concatenating $\hat{F}_{AV}$ and $\hat{F}_{VA}$ as $Z\in\mathbb{R}^{T_{l}\times 2D}$, MoDE performs cross-modal temporal aggregation and event dependency modeling on it. Specifically, MoDE consists of two branches: one for in-depth temporal modeling, and the other revealing multi-event relationships through a cascade of MoE layers. The outputs $Z_{t}\in\mathbb{R}^{T_{l}\times C}$ and $Z_{e}\in\mathbb{R}^{T_{l}\times C}$ from two branches are then added together to obtain $\hat{Z}\in\mathbb{R}^{T_{l}\times C}$ for final decoding. The details of MoDE are as shown in Section III-D.
@@
Decoder. The decoder $\mathscr{D}$, following [1, 5], includes a classification head and a regression head. Each of them consists of three layers of 1D convolutions and one layer activation function (Sigmoid for the classification head, ReLU for the regression head). The integrated features $\hat{Z}$ across all pyramid levels are fed into $\mathscr{D}$ to predict the probability $p(c_{n})$ of event categories and the distances $D_{e}=\{d^{s}_{n},d^{e}_{n}\}_{n=1}^{N}$ from the current moment $t$ to the temporal boundaries of events $T_{e}=\{t^{s}_{n},t^{e}_{n}\}_{n=1}^{N}$. Note that the regression head here is class-aware, its outputs are available only when audio-visual events are detected.
III-C Early Semantics Interaction
@@
Due to the lack of cross-modal semantic bridging in intermediate layers, the modality semantic gap lays obstacles to the understanding of audio-visual events in previous methods [1, 5]. To solve this issue, we propose Multi-Modal Early Fusion and Multi-Stage Semantic Guidance in ESI that enable hierarchical understanding of audio-visual events. The former enables the model to fuse audio-visual information early, while the latter performs progressive semantic bridging. The details are as shown in the lower left part of Figure 2:
@@
Multi-Modal Early Fusion. Before cross-modal semantic bridging, we need to fuse multi-modal information early so that the model can detect events that are both visible and audible among intermediate layers. As shown in Figure 2, our multi-modal early fusion can be divided into Single-Modal Attention, Audio(Visual)-Driven Mixture and Cross-Modal Pyramid, which are used for single-modal temporal modeling, multi-modal information fusion , and multiple temporal resolution modeling, respectively.
@@
(1) Single-Modal Attention. To capture long-term temporal relations among single-modal segments and filter out some noise, we perform single-modal temporal modeling as most methods. Specifically, we first optimize the original features through projection layers, where $F_{A}^{\prime}=f_{proj}(F_{A})$ and $F_{V}^{\prime}=f_{proj}(F_{V})$. Then the mapped audio and visual features $\{F_{A}^{\prime},F_{V}^{\prime}\}\in\mathbb{R}^{T\times D}$ are used for temporal aggregation through single-modal transformer blocks. Each single-modal transformer block includes a multi-headed self-attention (MSA) and a feed-forward network (FFN), which can be denote as:
@@
[…]
(2) Audio(Visual)-Driven Mixture. Our audio(visual)-driven mixture aims to enable audio(visual) features $\{\hat{F_{A}},\hat{F_{V}}\}$ to obtain complementary information from the audio-visual representations. We first need to get the early multi-modal representations, which contain synchronized (frame-aligned) multi-modal information. However, due to the significant differences between original audio and visual features $\{F_{A},F_{V}\}$, simply concatenating them may introduce much noise and lead to insufficient fusion. To solve this problem, we first design an aligner that includes two projection layers and a feature optimization attention block to map them onto the same representation space. The aligner is as follows:
@@
[…]
@@
[…]

### v2

@@ Abstract
Dense audio-visual event localization (DAVE) aims to identify event categories and locate the temporal boundaries in untrimmed videos. For such challenging task settings, most studies only employ audio-visual event semantic constraints on the final outputs, lacking progressive cross-modal semantic bridging in intermediate layers. This causes semantic gaps that hinder alignment between representations in audio and visual features, making it difficult to distinguish between event-related and irrelevant background content. Moreover, they rarely consider the correlations between events, which limits the model to infer co-occurring events among complex scenarios. In this paper, we incorporate multi-stage semantic guidance and multi-event relationship modeling, which respectively enable progressive semantic understanding of audio-visual events and adaptive extraction of event dependencies, thereby better focusing on event-related information. Specifically, our event-aware semantic guided network (ESG-Net) includes a early semantic interaction (ESI) module and a mixture of dependency experts (MoDE) module. ESI applys multi-stage semantic guidance to explicitly constrain the model in learning semantic information through multi-stage feature fusion and several classification loss functions, ensuring multi-stage understanding of event-related content. MoDE promotes the extraction of multi-event dependencies through multiple serial mixture of experts with adaptive weight allocation. Extensive experiments demonstrate that our method significantly surpasses the state-of-the-art methods, while greatly reducing parameters and computational load. Our code will be released on https://github.com/uchiha99999/ESG-Net.
@@ I Introduction
In recent years, most existing cross-modal frameworks for detecting dense events only perform event-related constraints (event classification and boundary localization losses) on the final outputs [1, 5]. While other methods [6] apply contrastive learning [7] for cross-modal alignment in intermediate layers at the feature level. However, the lack of semantic constraints during multi-modal feature fusion makes localizing audio-visual events in untrimmed videos particularly challenging. This is because the inherent temporal misalignment between audio and visual modalities creates significant semantic gaps (e.g., a visual scene of playing tennis accompanied by delayed cheers in Figure 1). Insufficient event supervision and forced alignment of multi-modal features without semantic constraints can introduce much noise due to mismatched video segments, making it difficult for the model to distinguish between event-related and irrelevant background content. Therefore, it is crucial to impose sufficient cross-modal semantic bridging in intermediate layers to achieve semantic alignment for accurate audio-visual event location.
@@
Figure 1: (a) Audio-visual events is intersection of visual and audio events, while significant differences exist between visual and audio content. (b) Most previous methods only performe semantic constraints on the final outputs. (c) We consider event-related semantic bridging between different modalities in intermediate layers through multi-stage fusion and semantic guidance.
@@
Additionally, few studies take the correlations between audio-visual events into consideration, which limits the model’s ability to infer co-occurring events in complex scenarios. Specifically, there are complex dependencies between different audio-visual events [1] (e.g., thunder is often accompanied by rain). These dependencies are variable (e.g., piano solo or playing together with guitar) and may differ in order and overlap duration. Approaches with fixed modeling structures are insufficient to fit such diverse dependency patterns. How to adaptively capture those dependencies in complex scenarios to improve the reasoning capability of the model for co-occurring events is also a crucial issue for DAVE.
@@
In this work, we propose an event-aware semantic guided network (ESG-Net), which leverages multi-stage semantic guidance for progressive audio-visual events understanding and multi-event relationship modeling for adaptive dependencies capture, respectively. (1) Inspired by audio-visual video parsing (AVVP) [8, 9, 10, 11], our model is designed to output event detection results from single-modal stage to multi-modal stages with consistent semantic supervision. As shown in Figure 1, compared with previous methods, ESG-Net can achieve cross-modal semantic bridging in intermediate layers through multi-stage constraints, effectively reducing the multi-modal semantic gap and guiding the model to focus on foreground information (e.g., audio-visual events like "playing tennis" tagged in Figure 1). (2) Additionally, we also consider the dependencies among multiple events and adaptively capture these complex multi-event relationships by mixture of experts (MoE) [12, 13, 14] module. Overall, our method progressively emphasizes event-related content through multi-stage processing, where the semantic bridging and dependency modeling jointly enable accurate audio-visual event detection.
@@
Specifically, ESG-Net consists of an early semantic interaction (ESI) module and a mixture of dependency experts (MoDE) module. (1) The ESI module includes multi-stage feature fusion and multi-stage semantic guidance. We progressively fuse multi-modal information and perform cross-modal semantic bridging by weighted classification loss functions across three stages: single-modal attention, audio(visual)-driven mixture, cross-modal pyramid. These stages respectively enable single-modal event detection, preliminary audio-visual event detection, and refined audio-visual event detection, ensuring sufficient semantic alignment between audio and visual branches in intermediate layers. (2) MoDE is proposed to address the the modeling of correlations between audio-visual events with efficient computation. MoDE consists of multiple sequential mixture of experts [12, 13, 14] layers, with each layer selecting only one expert at a time. Through flexible selection of experts in different layers, diverse event dependencies can be captured.
@@
Our event-aware semantic guided network can gradually focus on audio-visual content across multiple intermediate layers and capture multi-event relationships for accurate dense events localization. The results show that ESG-Net surpasses the baseline model by 2.1% on I3D+VGGish [15, 16] backbone and 4.1% on ONE-PEACE [17] backbone. Our contributions are summarized as:
@@
  • We introduce an early semantic interaction module and a mixture of dependency experts module, designed for cross-modal bridging by explicit semantic guidance in intermediate layers and adaptive multi-event dependencies capture, respectively.
  • Abundant experiments demonstrate that our event-aware semantic guided network outperforms state-of-the-art methods across various intersection-over-union thresholds.
@@ II-A Audio-Visual Event Localization
The audio-visual event localization (AVEL) task was first proposed by Tian et al. [18], who also introduced the AVE dataset and a dual multi-modal residual network. Then some studies [19, 20] provided supervised cross-modal fusion and event detection methods, while others [21, 22, 23, 24] are compatible with both weakly-supervised and supervised methods. Recently, Zhou et al. [25] proposed an open-vocabulary benchmark to address the closed-set limitation in AVEL, which enables recognition of unseen events. However, these studies are based on trimmed video clips, with a single event in each clip, which are not fit with the real-world scenarios. To address this issue, Geng et al. [1] released UnAV-100, a real-world dataset containing a large number of untrimmed videos with dense events, and introduced a benchmark for DAVE. While Krishna et al. [26] also proposed a similar task called dense video captioning (DVC), which simultaneously detects events and generates textual descriptions.
Localizing audio-visual events from untrimmed videos is a challenging setting, requiring models to simultaneously achieve multi-modal fusion and parse video segments. Vladimir et al. [27, 28] and Xie et al. [29] respectively achieved multi-modal DVC by integrating audio representations and globally-shared text representations with visual track. Geng et al. [5] subsequently proposed a unified framework to integrate three tasks, including temporal action localization (TAL) [30, 31, 32], sound event detection (SED) [33, 34], and AVEL [18, 19, 35]. Faegheh et al. [36] improved the weakly-supervised AVVP [8, 9, 10, 11] by employing a reference and an anchor module to extract single-modal and multi-modal information. Recently, Tang et al. [37] upgraded Vicuna to AVicuna, enhancing the multi-modal understanding capabilities of large language models by aligning audio and video features with text. Xing et al. [6] proposed a locality aware cross-modal correspondence learning framework and improved the inaccurate recognition of event boundaries [1] by using contrastive learning and dynamic local temporal attention. While Zhou et al. [38] improve the performance through adaptive cross-modal attention and temporal modeling at different granularities. However, these methods only employ event-related constraints on the final outputs, lacking cross-modal semantic bridging in intermediate layers, which remain modality semantic gap between audio and visual features. In this work, we perform multi-stage semantic guidance for multi-modal alignment, which allows the model to focus on event-related content in multiple intermediate layers.
@@ II-B Audio-Visual Bridging
The joint analysis of audio and video is a powerful tool that can be applied to various tasks [39]. How to reasonably bridge different modalities is crucial for models to understand task-related content. Early researches [40, 41, 42] simply concatenated and projected the audio and visual features extracted by sub-networks. This may result in insufficient features fusion and suboptimal extraction of task-specific information. Subsequently, researchers used the audio-visual similarity to guide attention toward relevant regions for particular tasks [18, 43, 44, 19, 45, 46]. Some methods also applied contrastive learning for self-supervised synchronous tasks [47, 48, 49]. Currently, there are typical frameworks mainly include single-modal modeling and multi-modal modeling, which are used for audio-visual correspondence (AVC) [50], audio-visual parsing (AVP) [51], and AVEL [22, 21, 10, 1, 52]. Most of these methods only empoly late multi-modal fusion and perform semantic constraints on the final outputs, ignoring semantic cross-modal bridging in intermediate layers, which may make the model difficult to focus on task-related content as early as possible.
Other studies attempt to use early cross-modal bridging to reduce modality gap. Introducing contrastive learning [7] is a reasonable method to strengthen audio-visual connections. For example, Hong et al. [53] proposed hyperbolic alignment to implement zero-shot learning. Trevine et al. [54] achieved video deepfake detection by aligning the multi-modal features of frames. Lee et al. [55] conducted pairwise contrastive learning to map audio, images, and text into a multi-modal shared latent space, while Mo et al. [56] built a framework for localization, separation, and recognition by matching positive samples. Although these methods can effectively optimize multi-modal features, there are many mismatched segments in most untrimmed videos. Forcibly performing frame-level alignment may introduce significant noise. In addition to contrastive learning, some studies used co-attention to capture the relationships between multi-modal inputs [57, 20, 23, 24, 1]. Other methods also added a parallel early fusion branch alongside two single-modal branches [58, 59, 60] for temporal alignment. However, these audio-visual bridging methods ignore the semantic alignment between different modalities in intermediate processes. Our method employs a multi-stage guided classification to enforce explicit cross-modal semantic bridging gradually, thereby reducing modality semantic gap and promoting the attention of our model to task-related content.
@@ III Method
To overcome the multi-modal semantic gap between audio and visual features in intermediate layers and the modeling of event correlations for DAVE, we design an early semantic interaction module to achieve multi-stage cross-modal semantic bridging, and a mixture of dependency experts module to capture complex relationships between events. An overview of the event-aware semantic guided network we propose is shown in Figure 2.
@@ III-B Overall Architecture
ESG-Net comprises four key components: 1) pre-trained audio encoder and visual encoder to extract features from original videos; 2) an early semantic interaction module for cross-modal semantic bridging; 3) a mixture of dependency experts module for capturing multi-event dependencies; 4) a decoder for localizing audio-visual events. These key components are executed in sequence as illustrated in Figure 2.
@@
Early Semantic Interaction Module. To achieve cross-modal semantic bridging in intermediate layers, ESI consists of Multi-Stage Feature Fusion and Multi-Stage Semantic Guidance. The Multi-Stage Feature Fusion includes: Single-Modal Attention for visual/audio modeling, Audio(Visual)-Driven Mixture for preliminary audio-visual modeling, and Cross-Modal Pyramid Module[1] for refined audio-visual modeling. The outputs of these three parts are $\{\hat{F}_{A},\hat{F}_{V}\}\in\mathbb{R}^{T_{m}\times D}$, $\{F_{AV},F_{VA}\}\in\mathbb{R}^{T_{m}\times D}$, $\{\hat{F}_{AV},\hat{F}_{VA}\}\in\mathbb{R}^{T_{l}\times D}$, respectively, where $T_{m}$ is the maximum length among all videos, $T_{l}$ is the total length of the multi-level pyramid features. Then, they are used for Multi-Stage Semantic Guidance through classfication loss functions. The details are as shown in Section III-C.
@@
Mixture of Dependency Experts Module. After concatenating $\hat{F}_{AV}$ and $\hat{F}_{VA}$ as $Z\in\mathbb{R}^{T_{l}\times 2D}$, MoDE performs cross-modal temporal aggregation and event dependency modeling. Specifically, MoDE consists of two branches: one for in-depth temporal modeling, and the other revealing multi-event relationships through a cascade of MoE layers. The outputs $Z_{t}\in\mathbb{R}^{T_{l}\times C}$ and $Z_{e}\in\mathbb{R}^{T_{l}\times C}$ from two branches are then added together to obtain $\hat{Z}\in\mathbb{R}^{T_{l}\times C}$ for final decoding. The details of MoDE are as shown in Section III-D.
@@
Decoder. The decoder $\mathscr{D}$, following [1, 5], includes a classification head and a regression head. Each of them consists of three layers of 1D convolutions and one layer of activation function (Sigmoid for the classification head, ReLU for the regression head). The integrated features $\hat{Z}$ across all pyramid levels are fed into $\mathscr{D}$ to predict the probability $p(c_{n})$ of event categories and the distances $D_{e}=\{d^{s}_{n},d^{e}_{n}\}_{n=1}^{N}$ from the current moment $t$ to the temporal boundaries of events $T_{e}=\{t^{s}_{n},t^{e}_{n}\}_{n=1}^{N}$. Note that the regression head here is class-aware, its outputs are available only when audio-visual events are detected.
III-C Early Semantic Interaction
@@
Due to the lack of cross-modal semantic bridging in intermediate layers, the modality semantic gap caused by temporal misalignment between audio and video lays obstacles to the understanding of audio-visual events in previous methods [1, 5]. To solve this issue, we propose Multi-Stage Feature Fusion and Multi-Stage Semantic Guidance in ESI that enable progressive understanding of audio-visual events. The former enables the model to fuse audio-visual information gradually, while the latter performs semantic bridging for progressive event understanding. The details are as shown in the lower left part of Figure 2:
@@
Multi-Stage Feature Fusion. Before cross-modal semantic bridging, we need to fuse multi-modal information gradually so that the model can better reduce modality gaps among intermediate layers. As shown in Figure 2, our multi-stage feature fusion can be divided into Single-Modal Attention, Audio(Visual)-Driven Mixture and Cross-Modal Pyramid, which are used for single-modal temporal modeling, multi-modal information fusion , and multiple temporal resolution modeling, respectively.
@@
(1) Single-Modal Attention. To capture long-term temporal relations among single-modal segments and filter out some noise, we perform single-modal temporal modeling. Following this stage, both the audio and visual modalities are able to preliminarily identify the approximate locations of events. Specifically, we first optimize the original features through projection layers, where $F_{A}^{\prime}=f_{proj}(F_{A})$ and $F_{V}^{\prime}=f_{proj}(F_{V})$. Then the mapped audio and visual features $\{F_{A}^{\prime},F_{V}^{\prime}\}\in\mathbb{R}^{T\times D}$ are used for temporal modeling through single-modal transformer blocks. Each single-modal transformer block includes a multi-headed self-attention (MSA) and a feed-forward network (FFN), which can be denote as:
@@
where $\{\hat{F_{A}},\hat{F_{V}}\}\in\mathbb{R}^{T_{m}\times D}$ are the outputs of audio and visual features, respectively. $W_{q}^{A}$, $W_{k}^{A}$, $W_{v}^{A}$, $W_{q}^{V}$, $W_{k}^{V},W_{v}^{V}$$\in\mathbb{R}^{D\times D}$ are learnable parameters.
(2) Audio(Visual)-Driven Mixture. After single-modal attention, we attempt to coarsely fuse audio and visual information to achieve preliminary perception of audio-visual events. Compared with direct interaction with another modality, our audio(visual)-driven mixture enables single-modal features to acquire complementary information from early multi-modal representations with smaller modality gaps (because early multi-modal representations retain some correlations and synchronized temporal structures).
We first build the early multi-modal representations, which contain frame-aligned audio-visual information. However, due to the significant differences between original audio and visual features $\{F_{A},F_{V}\}$, simply concatenating them to get the early multi-modal representations may introduce much noise and lead to insufficient fusion. So we first design an aligner that includes two projection layers and a feature optimization attention block to map them onto the same representation space:
@@
where $F_{g}^{\prime}\in\mathbb{R}^{T_{m}\times D}$ represents the early multi-modal representation, $W_{q}^{g}$, $W_{k}^{g}$, $W_{v}^{g}$$\in\mathbb{R}^{T_{m}\times T_{m}}$. Then, like the Single-Modal Attention in Eq. (1) and Eq. (2), the early multi-modal representation $F_{g}^{\prime}$ is also employed for temporal modeling:
@@
where $\hat{F_{g}}\in\mathbb{R}^{T_{m}\times D}$, $W_{q}^{g^{\prime}}$, $W_{k}^{g^{\prime}},W_{v}^{g^{\prime}}$$\in\mathbb{R}^{D\times D}$. Now the early multi-modal representation $\hat{F_{g}}$ contains frame-aligned audio and visual information. We introduce an audio(visual)-driven mixture to extract multi-modal complementary information for $\{\hat{F_{A}},\hat{F_{V}}\}$. The audio(visual)-driven mixture includes a multi-headed cross-attention (MCA) and a FFN, as follows: